Find control limits for a letter c with process average equal to nine disconformities
Answer #1

The average number of defects, c̅ = 9

UCL = c̅ + 3.c̅ = 9 + 3*√9 = 18
LCL = c̅ - 3.c̅ = 9 - 3*√9 = 0

So, the required control limit is [0, 18]